OpenStudy (anonymous):

find the inverse of the following transform " F(s) = e^(-2s) / (s+1)^2

OpenStudy (anonymous):

any one who can answer this?

OpenStudy (anonymous):

so we have to solve for s \[f(s) = \frac{ e^{-2s} }{ (s+1)^2}\]

OpenStudy (anonymous):

this is the question right ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

and we have to find F(t)

OpenStudy (anonymous):

((s+1)^2) * f(s) = e^(-2s) introduce ln on both sides : ln[((s+1)^2) * f(s)] = ln [e^(-2s)] ln[(s+1)^2] + ln[f(s)] = -2s 2ln(s+1) +2s = -ln[f(s)] ln(s+1) + ln(e^s) = -ln[f(s)] / 2 but then we still cant isolate s..

OpenStudy (anonymous):

no we have to use the t shifting theorm here

OpenStudy (anonymous):

dont know about it .. sorry cant help you